Suppose that the consumers spend 80% of each additional dollar of income. In other words, marginal propensity to consume (c1) is 0.8. Assuming a hypothetical economy which is composed of households and firms, what is the value of multiplier? QUESTION 27 Assume that the marginal propensity to consume is 0.8. How much will the output increase as a result of a $100 increase in investment spending? O 400 O 500 O 100 O 50 QUESTION 28 Assuming that there is no govemment spending or trade, an economys aggregate demand is given by its domestic consumption C and investment I, AD-CI 1Y+L. In the economys goods market equilibrium this equals its output: AD-Y. Solving for Y this yields: Given this equation, which of the following statements is correct? The multiplier is given by 1-c1 OThe boost in the economys output is not the same, regardless of whether the aggregate demand shock comes from an
Answer #1

1) multiplier = 1/(1 - MPC) = 1/0.2 = 5

2) Increase in output = Increase in investment x multiplier = 5 x 100 = $ 500